Consider a bit streaming scenario for a video where the following values apply: – The buffer size is 1 MiB – The low watermark is set at 100KiB – The high watermark is set at 900KiB – The incoming data rate is 1Mbps – The video display rate is 300 Kbps

 Assume that the video is playing and that the buffer content has dropped to the low-water mark. The media player sets the controls for data input to begin again



1
Expert's answer
2021-03-12T14:55:02-0500

Data which to be streamed:

(high watermark)-(low watermark)"=900-100=800\ kb"

Let "x" be the time after the streaming begins. Then:

(buffer size)"x-"(video display rate)"x=" (Data which to be streamed)

"1000x-300x=800"

"x=8/7\ s"

So, it will take "8\/7" second to fill up to the high water mark.